# GPT4Scene: Understand 3D Scenes from Videos with Vision-Language Models

### v2

@@ Abstract
In recent years, 2D Vision-Language Models (VLMs) have made significant strides in image-text understanding tasks. However, their performance in 3D spatial comprehension, which is critical for embodied intelligence, remains limited. Recent advances have leveraged 3D point clouds and multi-view images as inputs, yielding promising results. However, we propose exploring a purely vision-based solution inspired by human perception, which merely relies on visual cues for 3D spatial understanding. This paper empirically investigates the limitations of VLMs in 3D spatial knowledge, revealing that their primary shortcoming lies in the lack of global-local correspondence between the scene and individual frames. To address this, we introduce GPT4Scene, a novel visual prompting paradigm in VLM training and inference that helps build the global-local relationship, significantly improving the 3D spatial understanding of indoor scenes. Specifically, GPT4Scene constructs a 3D Bird’s Eye View (BEV) image from the video and marks consistent object IDs across both frames and the BEV image. The model then inputs the concatenated BEV image and video frames with markers. In zero-shot evaluations, GPT4Scene improves performance over closed-source VLMs like GPT-4o. Additionally, we prepare a processed video dataset consisting of 165K text annotation to fine-tune open-source VLMs, achieving state-of-the-art performance on all 3D understanding tasks. Surprisingly, after training with the GPT4Scene paradigm, VLMs consistently improve during inference, even without visual prompting and BEV image as explicit correspondence. It demonstrates that the proposed paradigm helps VLMs develop an intrinsic ability to understand 3D scenes, which paves the way for a noninvasive approach to extending pre-trained VLMs for 3D scene understanding.
@@ 1 Introduction
Embodied AI refers to intelligent systems capable of performing various tasks through interaction with the physical environment [41, 68, 10]. It has broad applications in industrial inspection, smart homes, and smart cities [41, 21, 10]. 3D scene understanding involves the ability of multi-modal language models to comprehend the overall layout of indoor environments and the spatial relationships between objects [6, 22, 13]. Thus, a strong foundation for embodied intelligence lies in understanding scenes effectively.
Currently, 3D point LLMs are a popular approach for understanding indoor scenes, using point clouds as input and aligning point features with LLMs to perform scene understanding tasks [21, 38]. However, this approach has several limitations. First, point clouds provide limited detailed information, such as fine geometric details, material properties, and intricate textures. Second, although some point LLMs [121, 21, 40] attempt to use point clouds and multi-images as inputs, they face challenges in aligning text, image, and point cloud modalities. Additionally, the significant imbalance in the data quantities of point clouds versus text/video data introduces further complications.
These limitations motivate us to explore indoor scene understanding using pure vision inputs. This approach aligns more closely with human perception, as people can understand 3D scenes without relying on explicit 3D data like point clouds. Vision Language Models (VLMs) have demonstrated impressive results in image-text multi-modal tasks[90, 62, 54, 119]. However, their application to understanding immersive 3D indoor scenes has not been well explored. We conduct a pilot study to investigate this potential by directly inputting a scene video into VLMs. Our results show that this approach causes VLMs to fail to understand 3D scenes. The core issue lies in the lack of global scene information and the misalignment between each frame’s local position and international context.
@@
To address this, we propose a framework called GPT4Scene to help VLMs establish spatial relationships as shown in Figure 1. First, we perform 3D reconstruction based on the input video to generate a Bird’s Eye View (BEV) image as additional input, providing a comprehensive perspective of the scene layout. We also incorporate Spatial-Temporal Object markers (STO markers) in both the 3D BEV image and 2D frames. These STO markers represent object IDs consistent across sequential frames (temporal level) and align with those in the 3D BEV image (spatial level), helping the VLMs establish the global-local relationship between the overall scene layout and the video frames.
In a zero-shot setting, we first apply GPT4Scene to powerful closed-source VLMs, such as GPT-4o, using video frames and the BEV image with STO markers. The performance is comparable to existing 3D point-based LLMs, demonstrating GPT4Scene’s effectiveness in enhancing VLMs’ 3D scene understanding capabilities. For smaller open-source VLMs, we construct a ScanAlign dataset, which includes video frames with STO markers, BEV images, and text annotations. After fine-tuning on this dataset, our method outperforms existing approaches. Moreover, after fine-tuning, VLMs exhibit impressive performance with only the raw scene video as input, whereas additional processing was necessary for good results before training. It indicates that GPT4Scene helps VLMs develop an intrinsic ability to model 3D scene relationships.
@@
  • We introduce two techniques: (1) A 3D BEV image with global context information and (2) Spatial-Temporal Object markers (STO markers) for spatial and temporal consistency across BEV image and video frames.
  • We construct ScanAlign, a dataset consisting of video frames, BEV images with STO markers, and text annotations. Fine-tuning VLMs on this dataset markedly enhances their 3D scene understanding abilities.
@@ 2 Related Work
3D indoor scene understanding. 3D indoor scene understanding allows robots to identify object positions, structures, and relationships within indoor environments, enabling question-and-answer interactions about the scene’s content. This process combines 3D perception with large language models (LLMs). 3D perception, as a foundational component, is typically trained on common indoor datasets [25, 8, 11, 69, 78, 89, 84, 28, 52, 116] using point clouds as input, supporting tasks like 3D object detection and instance segmentation [74, 48, 70, 82, 88, 96, 110]. Recent advancements in 3D Vision-Language Learning (3D-VL) combine 3D perception tasks with natural language, introducing diverse textual annotations on datasets like ScanNet to support tasks such as 3D Question Answering [6, 66, 106, 18], 3D Dense Captioning [22], and 3D Visual Grounding [13, 114, 2, 1, 51]. Initial studies focus on single 3D-VL tasks [108, 20, 97, 7, 36, 44, 46, 64, 104, 115, 33], while recent research introduc unified models for multiple tasks [9, 14]. The success of 2D-VL models is largely due to extensive 2D Vision-Language Pretraining (2D-VLP) with image-text pairs, leading some approaches to leverage 2D-VLP for 3D-VL [35, 37, 72, 86, 101, 112, 113, 99]. Though 3D-VLP has emerged [120, 121, 49, 29, 102, 47, 91], the limited availability of 3D point cloud-text pairs constrains its effectiveness. Therefore, achieving better results while utilizing 2D modality input would be beneficial.
@@
3D Point Cloud LLMs. 3D Vision Language tasks aim to integrate 3D scene understanding with natural language processing while incorporating LLMs, which enables more natural human-computer interaction. 3D point LLMs take point clouds as input, supporting natural language generation and interaction in 3D scenes. Early 3D LLMs focused on object-level geometry and appearance [34, 76, 100, 75]. Later, 3D LLMs expanded to indoor scenes, emphasizing spatial relationships among objects and overall scene features, often using cene point clouds with auxiliary 2D multi-view images [38, 21, 94, 32, 68]. To better capture object relationships, recent 3D LLMs decouple scene objects before inputting them into LLMs [41, 40]. However, they still lack detailed information about object materials and precise geometric details. Additionally, they miss the current first-person perspective, whereas existing embodied AI can determine the current position within a room based on the current frame [10, 118, 60]. Moreover, due to the limited amount of point cloud data, its general understanding capabilities are relatively limited compared to current VLMs. Therefore, using the video modality to understand 3D scenes has more practical value.
@@
Vision Language Models (VLMs). Vision Language Models (VLMs) are multimodal models that integrate visual and language processing capabilities, enabling the understanding and generation of combined image-text information. The origin of VLMs can be traced back to CLIP’s 2D image-text pair pretraining [77, 12, 81], which laid the foundation for incorporating LLMs. Early VLMs used attention mechanisms or Q-Former to fuse image and text modalities before inputting them to LLMs [55, 56, 27, 3, 45]. Later, an approach emerged that directly projects image features into the LLM using an MLP [61, 62, 54, 119, 90], achieving better performance. Building on this, VLMs expanded into visual grounding tasks [92, 73, 79, 53, 30, 85, 16] and further to video understanding by using spatiotemporal compression to process information from long image sequences [111, 24, 67, 59, 65, 4, 58, 83, 80, 5, 63]. However, VLMs cannot directly understand 3D scenes, so our task is to assist them in interpreting the 3D world.
3 Pilot Study
Figure 2: Pilot Study. We tested zero-shot 3D question answering on open-sourced VLM (Qwen2-VL) and closed-sourced VLM (GPT-4o). For GPT-4o, adding BEV images and corresponding object markers alongside video frames improved performance, but no significant improvement was seen with Qwen2-VL. While providing global features and connections between global and local information is helpful, smaller VLMs may need finetuning to enhance their 3D understanding capabilities. ‘VID’ means indoor video frames, ‘BEV’ means indoor BEV Map, and ‘Mrks’ means object ID markers on video frames and BEV maps.
Figure 3: The Framework of GPT4Scene. A scene video is processed by sampling frames, reconstructing a point cloud, and generating a BEV image. Object locations are detected from the point cloud and projected onto the video frames. The resulting frames and BEV image, enhanced with STO-Markers, are inputs for VLM training and inference.
In this paper, we opt for purely vision-based VLMs over 3D point LLMs. One reason is that 3D point clouds lack significant fine details. Another reason is that incorporating an additional 3D point cloud feature representation into pre-trained VLMs requires finetuning extensive point cloud data, which risks disrupting the well-established text-image alignment and potentially impairing the vision-language understanding capability. Additionally, accepting pure first-person perspective video input aligns more closely with the nature of human perception. Therefore, we choose VLMs to understand 3D scenes. The next question is: can VLMs effectively understand the 3D world directly?
Can VLMs understand the 3D world directly? To explore this issue, we conducted a pilot study, and the results are shown in Figure 2. We input a preprocessed scene video into the VLMs for the 3D question answering task. We find that neither the open-source Qwen2-VL-7B nor the powerful closed-source GPT-4o could perform the task effectively. We find that VLMs perform well in answering specific details about the scene but struggle with understanding the spatial relationships between objects. Therefore, we aim to provide VLMs with global information. We add the BEV image rendered from the indoor scene reconstructed from the input video. We find that this approach led to significant improvements in GPT-4o’s performance. Inspired by SOM [103], which enhances VLMs’ understanding by marking objects on images, we further explore whether providing markers on the video frames could improve VLMs’ spatial learning. We label the objects in the scene, marking them on both the video frames and the BEV image. The markers between the video frames and those on the BEV image are aligned. Inputting these labeled frames and BEV images into the VLMs led to further improvements in GPT-4o. Unfortunately, these two steps do not lead to improvements for Qwen2-VL. We believe establishing the connection between video frames and global information is useful. However, due to smaller-scale VLMs’ limited visual understanding capabilities, they cannot directly improve their 3D scene comprehension through zero-shot prompts alone. We may need to build a specific dataset and further finetune Qwen2-VL, which could lead to improved performance.
@@ 4 Methodology
In this section, we introduce our method. We present the GPT4Scene framework in Section 4.1, which enhances VLMs’ 3D scene understanding using video inputs. In Section 4.2, we discuss how we leverage zero-shot prompts to unlock the potential of powerful closed-source VLMs. In Section 4.3, we apply fine-tuning to enhance smaller-sized open-source VLMs for 3D understanding.
@@ 4.1 GPT4Scene Framework
Figure 3 shows GPT4Scene framework and explain how providing global scene information and object-level annotations improves the VLM’s understanding of 3D scenes. First, we assume the captured video is taken while moving around the indoor scene. It consists of $N$ frames, represented as $\bm{V}=\{I_{1},I_{2},\ldots,I_{N}\}$. Processing image sequences with VLMs poses challenges like limited image capacity, rapid context consumption, and higher inference costs. So we uniformly sample $n$ frames:
$s_{i}=[\ (i-1)\cdot N/n\ ]+1,\quad i=1,2,\ldots,n$ (1)
Here, $\{I_{s_{1}},I_{s_{2}},\ldots,I_{s_{n}}\}$ represents the sampled frames with indices $s_{i}$. We denote the sampled video as:
$\bm{V}_{\bm{s}}=\{I_{s_{1}},I_{s_{2}},\ldots,I_{s_{n}}\}$ (2)
This pre-selection dramatically reduces the time and cost of VLMs during training and inference without losing significant indoor scene information.
Global Information: 3D BEV Map. Egocentric video captures only local information, missing the broader scene context. To address this, we reconstruct the entire scene as a point cloud and render a panoramic image as a bird’s-eye view, giving the VLM a complete scene overview. Specifically, starting with an indoor scene video $\bm{V}=\{I_{1},I_{2},\ldots,I_{N}\}$ and the corresponding camera extrinsic $\bm{E}=\{E_{1},E_{2},\ldots,E_{N}\}$, we use 3D reconstruction techniques to generate 3D meshes and point clouds:
$\bm{P}=\mathcal{R}\left(\left\{(I_{n},E_{n})\mid n=1,2,\ldots,N\right\}\right)$ (3)
@@
$\bm{I}_{\bm{b}}=\mathcal{T}(\bm{P},E_{t})$ (4)
Here, $E^{t}$ denotes the camera extrinsic for the top-down view, and $\mathcal{T}(\cdot)$ represents the process of rendering the corresponding view based on the camera extrinsic, resulting in a BEV image of the scene. Notably, we continue to provide global 3D information to VLMs in the form of images.
Spatial-Temporal Object Markers. To help VLMs focus on specific objects, we introduce Spatial-Temporal Object markers (STO markers), ensuring consistency between 2D frames and the 3D BEV image. Obtain the 3D point cloud $\bm{P}$ reconstructed from the input video $\bm{V}$. Applying 3D instance segmentation like Mask3D [82] yields instance masks $\{M_{1},M_{2},\ldots M_{K}\}$, where $K$ denotes the total number of objects in the scene.
For BEV images, we first project the 3D mask onto the $xy$-plane and then extract the center coordinates of the bounding box formed by the projection, represented as $(C^{xy}_{1},C^{xy}_{2},\ldots,C^{xy}_{K})$, which are then displayed on the BEV image. For egocentric 2D markers, we first project $\{M_{1},M_{2},\ldots,M_{K}\}$ onto the video frames and then use the center of the bounding box formed by the 2D masks as the 2D markers, represented as $(C^{uv}_{1},C^{uv}_{2},\ldots,C^{uv}_{K})$. The 2D frames with markers and the BEV image are denoted as:
$\bm{V}_{\bm{s}}^{\prime}$ $=\bm{V}_{\bm{s}}(C^{uv}_{1},C^{uv}_{2},\ldots,C^{uv}_{K}),$ (5)
$\bm{I}_{\bm{b}}^{\prime}$ $=\bm{I}_{\bm{b}}(C^{xy}_{1},C^{xy}_{2},\ldots,C^{xy}_{K})$ (6)
Here, $\bm{V}_{\bm{s}}^{\prime}$ and $\bm{I}_{\bm{b}}^{\prime}$ refer to the videos or images with STO markers. We can observe that the 2D and 3D markers are spatially aligned, representing the same object. Additionally, $(C^{uv}_{1},C^{uv}_{2},\ldots,C^{uv}_{K})$ are also consistent across different frames, achieving temporal alignment.

### v4

@@ Abstract
In recent years, 2D Vision-Language Models (VLMs) have made significant strides in image-text understanding tasks. However, their performance in 3D spatial comprehension, which is critical for embodied intelligence, remains limited. Recent advances have leveraged 3D point clouds and multi-view images as inputs, yielding promising results. However, we propose exploring a purely vision-based solution inspired by human perception, which merely relies on visual cues for 3D spatial understanding. This paper empirically investigates the limitations of VLMs in 3D spatial knowledge, revealing that their primary shortcoming lies in the lack of global-local correspondence between the scene and individual frames. To address this, we introduce GPT4Scene, a novel visual prompting paradigm in VLM training and inference that helps build the global-local relationship, significantly improving the 3D spatial understanding of indoor scenes. Specifically, GPT4Scene constructs a Bird’s Eye View (BEV) image from the video and marks consistent object IDs across both frames and the BEV image. The model then inputs the concatenated BEV image and video frames with markers. In zero-shot evaluations, GPT4Scene improves performance over closed-source VLMs like GPT-4o. Additionally, we prepare a processed video dataset consisting of 165K text annotation to fine-tune open-source VLMs, achieving state-of-the-art performance on all 3D understanding tasks. Surprisingly, after training with the GPT4Scene paradigm, VLMs consistently improve during inference, even without object marker prompting and BEV image as explicit correspondence. It demonstrates that the proposed paradigm helps VLMs develop an intrinsic ability to understand 3D scenes, which paves the way for a seamless approach to extending pre-trained VLMs for 3D scene understanding.
@@ 1 Introduction
3D scene understanding aims to comprehend the overall layout of the surrounding complex environments and the spatial relationships between objects [5, 21, 12]. It plays a crucial role in applications such as embodied intelligence, virtual reality, and smart cities [40, 20, 9]. With the rapid development of Large Language Models [91, 89, 71, 92, 90], Vision-Language Models have demonstrated impressive performance in image and video understanding [60, 61, 53, 126, 96]. Researchers have extended this paradigm to 3D perception by incorporating point clouds, aiming to improve scene understanding [37, 20, 100, 31, 68, 40, 39, 49].
Recent 3D point LLMs leverage point clouds aligned with LLM features for indoor scene understanding [20, 37]. While combining point clouds and images [128, 20, 39] in Point-Vision-LLM paradigms improves performance through richer visual cues, aligning these modalities with text remains challenging. This complexity motivates exploration of vision-only solutions, inspired by humans’ natural ability to perceive 3D environments without point clouds, offering a promising direction for efficient scene analysis.
In this work, we aim to leverage pre-trained VLMs without modifying their architecture, maximizing their vision perception capabilities. However, their effectiveness in understanding immersive 3D indoor scenes remains limited. Our analysis shows that directly inputting scene videos into VLMs fails in 3D scene understanding due to two factors: i) the lack of a global scene representation, ii) misalignment between per-frame local observations and their spatial-temporal context.
@@
To address this, we propose GPT4Scene, a framework that enhances VLMs’ spatial understanding (see Figure 1). We first perform 3D reconstruction on input videos to generate a Bird’s Eye View (BEV) image, offering a comprehensive scene layout. Additionally, we introduce Spatial-Temporal Object markers (STO-markers) in both the BEV image and 2D frames. These markers maintain consistent object IDs across frames (temporal level) and align with the BEV image (spatial level), bridging the global-local relationship. Empirical results show that GPT4Scene remains robust to reconstruction quality and marker accuracy, as it prioritizes learning global-local correspondences over precise geometric reconstructions.
We first explored the effectiveness of GPT4Scene under a training-free approach. Experimental results revealed that it was notably effective for powerful large-scale VLMs such as Qwen2-VL-72B [96], as well as closed-source models like GPT-4o [72] and Gemini-1.5-Pro [88], significantly enhancing their 3D scene understanding capabilities, even reaching levels comparable to previous state-of-the-art point-based methods. However, the improvements were limited on smaller-scale models. For smaller open-source vision-language models (VLMs), we introduce ScanAlign, a multimodal dataset comprising 165K aligned data pairs featuring STO-marker-annotated video frames, BEV images, and textual descriptions. Fine-tuning Qwen2-VL-7B on ScanAlign achieves state-of-the-art performance, delivering a 48% relative improvement (40.7 → 60.7 EM-1 score) on 3D question answering (SQA3D), surpassing the previous SOTA Chat-Scene by 11.0% (54.6 vs 60.7). The model demonstrates even stronger 3D visual grounding capabilities on Multi3DRef, outperforming Chat-Scene by 13.0% (57.1 → 64.5). These advancements underscore GPT4Scene’s effectiveness in seamlessly enhancing VLMs with strong 3D spatial understanding.
@@
  • We introduce two techniques: i) A 3D BEV image with global context information, and ii) Spatial-Temporal Object markers (STO-markers) for spatial and temporal consistency across BEV image and video frames.
  • We introduce ScanAlign, a dataset comprising video frames, BEV images with STO-markers, and text annotations. Fine-tuning VLMs on this dataset significantly improves 3D understanding, even with raw video inputs.
@@ 2 Related Work
3D Indoor Scene Understanding. 3D indoor scene understanding allows robots to identify object positions, structures, and relationships within indoor environments, enabling question-and-answer interactions about the scene’s content. This process combines 3D perception with large language models (LLMs). 3D perception, as a foundational component, is typically trained on common indoor datasets [24, 7, 10, 69, 79, 95, 85, 27, 51, 124] using point clouds as input, supporting tasks like 3D object detection and instance segmentation [75, 47, 70, 83, 94, 102, 116]. Recent advancements in 3D Vision-Language Learning (3D-VL) combine 3D perception tasks with natural language, introducing diverse textual annotations on datasets like ScanNet to support tasks such as 3D Question Answering [5, 66, 112, 17], 3D Dense Captioning [21], and 3D Visual Grounding [12, 121, 2, 1, 50]. Initial studies focus on single 3D-VL tasks [114, 19, 103, 6, 35, 43, 45, 64, 110, 122, 32], while recent research introduces unified models for multiple tasks [8, 13]. 2D vision-language pretraining (2D-VLP) has driven progress in 3D visual-language learning (3D-VL) [34, 36, 73, 87, 107, 118, 119, 105], with recent 3D-VLP [127, 128, 48, 28, 108, 46, 97] methods demonstrating that combining 2D visual cues with 3D point clouds enables complementary cross-modal alignment. This multimodal approach helps overcome geometric complexity and sparse annotation challenges in pure point cloud processing.
@@
3D Point Cloud LLMs. 3D vision-language tasks aim to integrate 3D scene understanding with natural language processing. However, we aspire to go further by incorporating 3D content into large language models (LLMs) to achieve more natural human-computer interaction. Initially, this began with 3D point cloud LLMs. 3D point cloud LLMs take point clouds as input, enabling natural language generation and interaction in 3D scenes. Early 3D LLMs focused on object-level geometry and appearance [33, 77, 106, 76]. Later, they expanded to indoor scenes, emphasizing spatial relationships among objects and overall scene features, often utilizing scene point clouds augmented with auxiliary 2D multi-view images [37, 20, 100, 31, 68]. To better capture object relationships, recent 3D LLMs decouple scene objects before feeding them into LLMs [40, 39]. Some approaches rely more heavily on visual inputs to determine scene context [9, 125, 59]. Here, we aim to explore whether pure visual inputs can better handle indoor scene understanding.
@@
Vision Language Models (VLMs). Vision Language Models (VLMs) are multimodal models that integrate visual and language processing capabilities, enabling the understanding and generation of combined image-text information. The origin of VLMs can be traced back to CLIP’s 2D image-text pair pretraining [78, 11, 82], which laid the foundation for incorporating LLMs. Early VLMs used attention mechanisms or Q-Former to fuse image and text modalities before inputting them to LLMs [55, 56, 26, 3, 44]. Later, an approach emerged that directly projects image features into the LLM using an MLP [60, 61, 53, 126, 96], achieving better performance. Building on this, VLMs expanded into visual grounding tasks [98, 74, 80, 52, 29, 86, 15] and further to video understanding by using spatiotemporal compression to process information from long image sequences [117, 23, 67, 65, 4, 58, 84, 81, 62]. Currently, while some studies employ VLMs for indoor scene understanding, most remain either benchmark-oriented [109, 54] or still rely on 3D-based methods [120, 123]. This fundamentally reveals that VLMs cannot directly comprehend 3D scenes, making our core mission to empower them with 3D world interpretation capabilities.
@@ 3 Methodology
Figure 2: The Framework of GPT4Scene. A scene video is processed by sampling frames, reconstructing a point cloud, and generating a BEV image. Object locations are detected from the point cloud and projected onto the video frames. The resulting frames and BEV image, enhanced with STO-markers, are inputs for Large Language Model (VLM) training and inference.
In this section, we introduce the GPT4Scene framework as shown in Section 3.1, which enhances VLMs’ 3D scene understanding using video inputs. In Section 3.2, we discuss how we leverage zero-shot prompts to unlock the potential of powerful large-scale VLMs. In Section 3.3, we apply fine-tuning to enhance VLMs for 3D understanding.
@@ 3.1 GPT4Scene Framework
This section first presents an overview of the GPT4Scene framework, then details the 3D BEV Map (providing global scene layouts) and STO-markers (delivering localized object cues) as core components.
Overview. Here we introduce the architecture of GPT4Scene. Figure 2 shows how GPT4Scene integrates global scene layouts and object-level details to improve VLMs’ 3D scene understanding. Given a video $\mathcal{V}\!=\!\{I_{1},\ldots,I_{N}\}$ captured by moving around an indoor scene, we first approximately uniformly sample $n$ frames using indices: $s_{i}=\left\lfloor(i-1)\frac{N}{n}\right\rfloor+1,\quad\forall i\in\{1,\ldots,n\}$ to form $\mathcal{V}^{*}\!=\!\{I_{s_{1}},\ldots,I_{s_{n}}\}$. This strategy reduces token counts and computational overhead for Vision-Language Models (VLMs) while preserving scene context, without significant information loss. We then leverage the complete temporal sequence for 3D scene reconstruction, generating a holistic Bird’s-Eye View (BEV) map. Subsequent 3D instance segmentation enables precise object localization, which is projected onto both the BEV map and 2D video frames to establish Spatio-Temporal Object markers (STO-markers). The technical details are elaborated as follows.
Global Information: 3D BEV Map. Egocentric videos lack global scene context. We address this by reconstructing the 3D scene into a point cloud and rendering a bird’s-eye view (BEV) image for holistic VLM understanding. Given video $\mathcal{V}=\{I_{1},I_{2},\ldots,I_{N}\}$ with camera extrinsics =$\mathcal{E}=\{E_{1},E_{2},\ldots,E_{N}\}$, we use 3D reconstruction techniques to generate 3D meshes and point clouds:
$\mathcal{P}=\mathcal{R}\left(\{(I_{t},E_{t})\}_{t=1}^{N}\right)$ (1)
@@
$\mathcal{I}_{b}=\mathcal{T}(\mathcal{P},\ E_{top})$ (2)
Here, $E_{\text{top}}\in\text{SE}(3)$ denotes the camera extrinsic for the top-down view, and $\mathcal{T}(\cdot)$ represents the process of rendering the corresponding view based on the camera extrinsic, resulting in a BEV image of the scene. In particular, we continue to provide global 3D information to VLMs in the form of the top-view images instead of points.
Local Correspondence: STO-markers. To help VLMs focus on specific objects, we introduce Spatial-Temporal Object markers (STO-markers), ensuring consistency between 2D frames and the 3D BEV image. Obtain the 3D point cloud $\mathcal{P}$ reconstructed from the input video $\mathcal{V}$. Applying 3D instance segmentation like Mask3D [83] yields instance masks =$\mathcal{M}=\{M_{1},M_{2},\ldots M_{K}\}$, where $K$ denotes the total number of objects in the scene.
For the BEV image $\mathcal{I}_{\bm{b}}$, we begin by projecting the 3D segmentation masks $\mathcal{M}$ onto the $xy$ plane. Then we extract the center coordinates of the resulting bounding box, denoted as $\bm{C}^{\bm{x}\bm{y}}=\{C^{xy}_{1},C^{xy}_{2},\ldots,C^{xy}_{K}\}$. These coordinates are subsequently overlaid onto the BEV image.
For selected egocentric video frames $\mathcal{V}^{*}$, the 3D instance masks $\mathcal{M}$ are projected onto each frame according to its corresponding camera pose. For each frame $i$, we extract the centroid of the 2D mask of each object as its 2D marker, defined as $\bm{C}^{\bm{u}\bm{v}}_{i}=\{C^{uv}_{i,1},C^{uv}_{i,2},\ldots,C^{uv}_{i,K}\}$ , where $C^{uv}_{i,k}$ represents the 2D marker for the $k$-th object in the $i$-th frame. The processed 2D frames with markers and the BEV image with markers are defined as follows:
$\mathcal{V}^{*\prime}$ =$=\left\{\mathcal{F}\left(I_{i},\ \bm{C}_{i}^{\bm{u}\bm{v}}\right)\mid i=s_{1},s_{2},\dots,s_{n}\right\}$ (3)
$\mathcal{I}_{b}^{\prime}$ $=\mathcal{F}\left(\mathcal{I}_{\bm{b}},\ \bm{C}^{\bm{x}\bm{y}}\right)$ (4)
Here, $\mathcal{F}(\cdot)$ represents the operation of overlaying markers onto images, $\mathcal{V}^{*\prime}$ and $\mathcal{I}_{b}^{\prime}$ denote the video frames and BEV image with STO-markers. Notably, the 2D markers in each frame and their corresponding 3D markers in the BEV image maintain spatial alignment, ensuring that they represent the same objects. Moreover, $\bm{C}^{\bm{u}\bm{v}}_{i}$ remains consistent across frames at the object level, preserving temporal coherence.